a projectile is projected from the ground has its direction of motion making an angle of 45 degrees with the horizontal at a height of 40m. its initial velocity of projection is 50m/s,the angle of projection is 

Grade:12

3 Answers

akash
8 years ago
direction of motion making 45 degrees means at that point its horizontal and vertical velocity are same. now try to solve.
Dayne Fernandes
8 years ago
solved!
Mansi arya
4 years ago
Let say angle of projection = z
Velocity is 50m/s
 
Vertical speed at time of projection= 50sinz
 
Horizontal speed at time of projection= 50cosz
Angle at 40m is 45°
 
Let say velocity is V at 40m height
50cosz = Vcos45
= V =50√2 cosz
Vsin45 is vertical velocity = 50cosz
Using formula v² - u² = 2as 
V = Vsin45 = 50cosz 
U= 50sinz
a =-10
50²( cos²z- sin²z )= -800
Cos² z- sin²z = -8/25
Cos2z = -8/25
2z= cos-¹ (-8/25)
Z= 1/2 cos-¹(-8/25)
 
Hope you find it useful
S =40
50²()